### v1

# Channel Coding for Unequal Error Protection in Digital Semantic Communication

Abstract
Semantic communication is an emerging paradigm that prioritizes transmitting task-relevant information over accurately delivering raw data bits. In this paper, we address an unequal error protection (UEP) problem in digital semantic communication, where bits of higher semantic importance require stronger protection. To quantify bit-level importance, we leverage bit-flip probabilities of semantic bits as target error protection levels, which are jointly learned with semantic encoder and decoder. We propose two novel channel coding frameworks aimed at minimizing the total blocklength while satisfying UEP constraints. First, we develop a bit-level UEP framework based on repetition coding, in which the repetition number for each bit is optimized to precisely meet its target bit-flip probability. Second, we introduce a block-level UEP framework utilizing modern channel codes, where semantic bits with similar target bit-flip probabilities are grouped to exploit coding gains. Within this framework, we propose a bit-grouping algorithm guided by finite blocklength capacity analysis. Simulation results conducted on image transmission tasks confirm that the proposed frameworks significantly outperform conventional approaches, yielding substantial improvements in both task performance and transmission efficiency.
Index Terms:
Semantic communication, Semantic importance, Unequal error protection, Channel coding, Repetition coding
I Introduction
Semantic communication is an emerging and disruptive paradigm that prioritizes the transmission of task-relevant information, rather than the accurate delivery of raw data bits as in traditional communication systems [1, 2, 3, 4, 5]. By enabling task-aware transmission and reception, semantic communication facilitates more efficient utilization of communication resources and improves robustness to channel impairments such as fading, interference, and noise. In the early stages of research, most efforts focused on extracting abstract semantic features from source data while accounting for channel conditions and task performance.
A representative direction in this context is neural network-based joint source-channel coding (JSCC), which integrates source and channel coding/decoding through a neural-type encoder/decoder architecture. Typically, the semantic encoder and decoder, placed at the transmitter and receiver respectively, are jointly trained in an end-to-end manner to minimize a task-specific loss function. Neural JSCC approaches have demonstrated substantial performance gains over traditional separate source and channel coding approaches, especially in scenarios with low signal-to-noise ratio (SNR) or limited bandwidth [6, 7, 8, 9]. However, since neural encoders generally produce continuous-valued outputs, these approaches often rely on analog transmission, which lacks compatibility with modern digital communication systems and suffers from limited flexibility in rate control.
Digital semantic communication has gained increasing attention due to its superior compatibility and flexibility compared to the analog JSCC approach [10, 11, 12, 13, 14]. A representative work is [10], where the end-to-end training of the semantic encoder and decoder was performed under finite constellation constraints (e.g., quadrature amplitude modulation), with constellation learning enabled via a differentiable soft-to-hard quantization layer. In [11, 12], digital outputs were implemented by quantization of neural encoder outputs. To enable backpropagation through the quantization process, they utilized the Gumbel-softmax trick for text transmission and multi-task scenarios, respectively. For a modulation-free approach, in [13], semantic-related bits were generated by interpreting the neural encoder outputs as sampling probabilities, while differentiability was preserved using a low-variance gradient estimator. In [14], digital symbols were produced by sampling from encoder outputs, followed by digital modulation. A channel-adaptive modulation scheme was also introduced in [14] to ensure robust task performance across varying SNRs. However, none of these methods explicitly quantify the importance of semantic features, limiting their ability to integrate with diverse channel coding and modulation schemes.
Recent studies have attempted to quantify the importance of various types of semantic features in semantic communication [15, 16, 17, 18]. In [15], the inherent importance imbalance introduced by multi-level quantization was explored. Since quantization errors vary depending on the bit positions, a bit-grouping strategy was developed for modulation, treating the most significant bit (MSB) after quantization as the most important. However, this approach considers quantization error as the only source of bit-level importance and does not account for semantic relevance. In [16, 17], the attention score of an image patch, extracted from a vision transformer, was used as a measure of semantic importance. Based on these scores, an importance-aware selection or quantization scheme was developed to enhance image transmission efficiency. However, such approaches are limited to image-related tasks due to their reliance on vision transformers. A universal approach was studied in [18], where the bit-flip probability of a binary symmetric channel (BSC) was interpreted as a measure of bit-level significance in semantic communication. Based on this interpretation, power allocation and modulation were optimized to meet the desired bit-flip probabilities. Despite these efforts, none of the existing methods consider channel coding design based on semantic importance, which is an essential component for ensuring reliable communication, particularly under constrained power and low SNR conditions.
Channel coding has long been a key ingredient for addressing the unequal error protection (UEP) problem in traditional communication systems [19, 20, 21]. For instance, in [19], rate-compatible punctured convolutional codes (RCPC) were proposed to assign higher code rates to less important information blocks. However, the exact error probability analysis of convolutional codes is intractable, making real-time rate adaptation difficult in practice. In [20], the assignment of code rates was optimized to different source blocks for progressive transmission using dynamic programming. While the method achieves optimality, it incurs excessive computational overhead, and estimating the expected distortion of each state remains a challenge, particularly in the context of semantic communication. In [21], it was shown that certain bit positions inherently possess stronger error-correction capabilities, and formal conditions were derived to characterize this behavior. Nonetheless, exploiting this potential requires maximum-likelihood decoding, thereby limiting the use of efficient bounded-distance decoding algorithms such as the Berlekamp-Massey algorithm [22].
Only a limited number of studies have investigated the UEP problem in the context of semantic communication [23, 24]. In [23], the authors addressed UEP for multi-modal semantic communication by applying the robust verification problem (RVP), which identifies an input feature as important if small perturbations lead to large deviations at the decoder output. In [24], the bits corresponding to segmentation maps were considered less important than those of latent maps, and the inherent UEP capability of polar codes was leveraged to provide stronger protection for the latent map bits. However, all these methods offer only modality-level or source-level error protection, rather than bit-level granularity. Addressing bit-level error protection is critical in digital semantic communication, as it enables fine-grained control over channel coding and provides greater potential for improving coding efficiency and resource utilization. To the best of the authors’ knowledge, no prior work has explored the bit-level UEP problem in digital semantic communication.
Figure 1: An illustration of a digital semantic communication system with the proposed UEP framework.
To fill this research gap, we present novel channel coding frameworks that address the bit-level UEP problem in digital semantic communication. In these frameworks, we extract the target error protection levels of semantic bits by leveraging their bit-flip probabilities, jointly learned with the semantic encoder and decoder. Based on the extracted levels, we formulate the bit-level UEP problem as the minimization of the total blocklength, subject to the constraint that each semantic bit’s actual bit-flip probability does not exceed its target value. To solve this problem, we propose a repetition coding-based framework that optimizes the number of repetitions assigned to each semantic bit. To further enhance coding gain, we also develop a block-level UEP framework utilizing modern channel codes such as polar codes and low-density parity-check (LDPC) codes. In this framework, we apply block coding to groups of semantic bits with similar target bit-flip probabilities. Simulation results demonstrate that the proposed frameworks significantly improve both task performance and transmission efficiency in digital semantic communication systems. The main contributions of this paper are summarized as follows:
  • We formulate a novel channel coding problem for digital semantic communication that explicitly incorporates the inherent UEP nature of semantic bits. The objective is to minimize total blocklength while ensuring that each bit’s flip probability does not exceed its respective target. Leveraging the learned bit-flip probabilities of semantic bits as target protection levels is a new perspective that has not been explored in the existing literature.
  • We propose a bit-level UEP framework based on repetition coding, in which the number of repetitions assigned to each semantic bit is optimized to satisfy its target bit-flip probability. To solve this efficiently, we develop a low-complexity bisection algorithm for determining the minimum required repetition count per bit. This is the first work to apply repetition coding for enforcing bit-level semantic importance in digital semantic communication.
  • We present a block-level UEP framework in which semantic bits with similar protection levels are grouped and encoded using polar codes or LDPC codes. To support this, we introduce a bit-grouping algorithm guided by finite blocklength capacity analysis. This framework is the first to address bit grouping for UEP from a finite blocklength perspective in semantic communication systems.
The remainder of the paper is organized as follows. In Sec. II, we describe the system model of a digital semantic communication scenario and introduce an end-to-end training method to extract bit-wise error protection levels. In Sec. III, we formulate a bit-level UEP problem and propose a repetition coding framework to solve this problem. In Sec. IV, we develop a block-level UEP framework that groups semantic bits with similar protection levels. In Sec. V, we provide simulation results that validate the effectiveness of the proposed UEP frameworks across various datasets. Finally, in Sec. VI, we present our conclusions and future research directions.
II System Model and Preliminary
In this section, we describe the digital semantic communication scenario considered in this work. We then introduce the end-to-end training method proposed in [18], which is employed to extract the target error protection levels of semantic bits.
II-A Digital Semantic Communication Scenario
In this work, we consider a digital semantic communication system for image transmission [^1] In this work, we focus on an image transmission task for simplicity; however, the core components of the proposed framework are not limited to this setting and can be readily extended to other types of semantic sources and tasks., as illustrated in Fig. 1. Given an image data $\bm{u}\in\mathbb{R}^{U}$, the transmitter extracts a semantic feature from the input data using a semantic encoder as follows:
$\bm{v}=f_{{\bm{\theta}}_{\text{enc}}}(\bm{u})\in\mathbb{R}^{M},$ (1)
where $f_{{\bm{\theta}}_{\text{enc}}}$ denotes the semantic encoder parameterized by $\bm{\theta}_{\rm enc}$, and $\bm{v}$ represents the semantic feature vector of length $M$. After encoding, each element of $\bm{v}$ is quantized using a uniform $B$-bit quantizer. The resulting quantized output is defined as
${\bm{q}}_{i}=\textsf{Q}(v_{i})\in\mathcal{Q},\quad i\in[M],$ (2)
where $v_{i}$ denotes the $i$-th element of $\bm{v}$, $\mathcal{Q}=\{\tilde{\bm{q}}_{1},\tilde{\bm{q}}_{2},\ldots,\tilde{\bm{q}}_{2^{B}}\}$ represents the quantizer codebook, and each $\tilde{\bm{q}}_{i}$ is a binary vector of length $B$. In this work, we refer to each binary entry of $\tilde{\bm{q}}_{i}$ as a semantic bit. The transmitter then constructs a semantic bit sequence as $\bm{b}=[\bm{q}_{1}^{\top},\cdots,\bm{q}_{M}^{\top}]^{\top}\in\{0,1\}^{K}$, where $K=MB$ is the total number of semantic bits. To ensure robust transmission of these bits, channel coding is applied, which augments redundant bits for error correction. The encoded bit sequence is represented as
$\tilde{\bm{b}}=g_{\rm enc}({\bm{b}})\in\{0,1\}^{N_{\rm tot}},$ (3)
where $g_{\rm enc}$ denotes the channel encoder. As a result of this process, the $K$ semantic bits are extended to $N_{\rm tot}$ coded bits. Details of the channel encoding process considered in our work will be presented in Sec. III and Sec. IV. This encoded bitstream $\tilde{\bm{b}}$ is subsequently mapped to a symbol sequence $\bm{x}=[x_{1},\cdots,x_{T}]^{\top}\in\mathcal{X}^{T}$ by digital modulation, where $\mathcal{X}$ is a constellation set and $T$ is the length of the symbol sequence.
We consider an additive white Gaussian noise (AWGN) channel to model the wireless link between the transmitter and the receiver. Under this channel model, the received signal at time slot $t$ is given by
$y_{t}=\sqrt{p_{t}}x_{t}+v_{t},\quad t\in[T],$ (4)
where $v_{t}\sim\mathcal{CN}(0,\sigma^{2})$ denotes Gaussian noise with zero mean and variance $\sigma^{2}$, and $p_{t}$ is the transmission power allocated to the $t$-th symbol.
Upon receiving the signal, the receiver attempts to recover a semantic bit sequence $\hat{\bm{b}}=[\hat{\bm{q}}_{1}^{\top},\cdots,\hat{\bm{q}}_{M}^{\top}]^{\top}\in\{0,1\}^{K}$ using channel decoding, as follows:
$\hat{\bm{b}}=g_{\rm dec}(\bm{y}),$ (5)
where $g_{\rm dec}$ is the channel decoder, and ${\bm{y}}=[y_{1},\cdots,y_{T}]^{\sf T}$ is a received signal vector. In our work, we assume that symbol demodulation is performed as a part of channel decoding. After reconstructing the semantic bit sequence, the receiver recovers each semantic feature through the dequantization process ${\sf Q}^{-1}$, as follows:
$\hat{v}_{i}={\sf Q}^{-1}(\hat{\bm{q}}_{i})\in\mathbb{R},\quad i\in[M].$ (6)
The recovered feature vector is then fed into the semantic decoder to reconstruct the original input data:
$\hat{\bm{u}}=f_{\bm{\theta}_{\rm dec}}(\hat{\bm{v}})\in\mathbb{R}^{U},$ (7)
where $f_{\bm{\theta}_{\rm dec}}$ denotes the semantic decoder parameterized by $\bm{\theta}_{\rm dec}$.
Figure 2: An illustration of the training procedure of the end-to-end training method in [18] along with the resulting bit-flip probabilities for the MNIST dataset when $\lambda=10^{-4}$.
II-B Error Protection Level Extraction via End-to-End Training
In this work, we adopt the end-to-end training method proposed in [18] to extract the target error protection levels of the semantic bits in $\bm{b}$}. In this method, the entire digital communication process, from the transmission of $\bm{b}$ at the transmitter to the reconstruction of $\hat{\bm{b}}$ at the receiver, is encapsulated as $K$ parallel BSCs, where the $i$-th channel has a bit-flip probability $\mu_{i}$. Then, the set of bit-flip probabilities, $\{{\mu_{i}}\}_{i=1}^{K}$, are jointly optimized with the semantic encoder and decoder based on the following loss function [18]:
[…]
where the first term minimizes the reconstruction error, and the second term penalizes the deviations of bit-flip probabilities from the maximum bit-flip value of 0.5 with regularization weight $\lambda>0$. To enable the training of $\mu_{i}$ in a differentiable manner, a continuous relaxation technique is employed in [18].
The end-to-end training method in [18] effectively learns the bit-flip probabilities of the semantic bits while accounting for the impact of each bit’s flip probability on task performance. Moreover, the semantic encoder learns a mapping that assigns important features to semantic bits associated with lower bit-flip probabilities. In other words, this framework jointly optimizes the feature-to-bit mapping and the corresponding bit-flip probability values. Fig. 2 illustrates the training procedure of this method, along with the resulting bit-flip probabilities for the MNIST dataset when $\lambda=10^{-4}$. Inspired by the ability of this method to capture bit-level importance, we adopt the learned bit-flip probabilities as target error protection levels in our framework.
Numerical Example (Validation of Unequal Error Protection Levels): To verify the validity of the unequal error protection levels extracted using the method in [18], we present a numerical example that demonstrates the impact of protecting different semantic bits on task performance. In this example, the semantic bit sequence of a test image is sorted in ascending order of the learned bit-flip probabilities. The sorted sequence is then divided into eight segments, with the first segment containing the most critical bits. To evaluate semantic robustness, we flip the bits in each segment individually while keeping the remaining segments unchanged. Fig. 3 visualizes reconstructed CIFAR-10 images and the corresponding peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) values when the $i$-th segment is flipped. As $i$ increases, the bits tend to have higher bit-flip probabilities, indicating lower semantic importance for the downstream task. Accordingly, both PSNR and SSIM improve with increasing $i$. Notably, flipping the first segment causes a substantial degradation in reconstruction quality, even resulting in a negative SSIM value, whereas flipping the eighth segment has minimal impact on the image. These results confirm that the semantic encoder learns to associate semantic importance with varying bit-flip probabilities, effectively translating them into the required levels of error protection.
Figure 3: Visualization of reconstructed CIFAR-10 images and corresponding PSNR and SSIM values when flipping the $i$-th segment of sorted semantic bits.
III Channel Coding for Bit-Level UEP
in Semantic Communication
In this section, we formulate the bit-level UEP problem for semantic communication, in which the learned bit-flip probabilities are set as target error protection levels. We then present a channel coding method to solve this problem based on repetition coding.
III-A Problem Formulation for Bit-Level UEP
In the bit-level UEP framework, our goal is to protect each semantic bit so that its actual bit-flip probability does not exceed the target bit-flip probability $\mu_{i}$. To this end, we formulate a bit-level UEP problem, aiming to minimize the total blocklength while satisfying the target bit-flip probability constraints for the semantic bits. Let ${\sf BER}(\{i\},n_{i})$ denote the bit error probability of the $i$-th semantic bit when it is encoded into $n_{i}$ bits. Note that ${\sf BER}(\{i\},n_{i})$ depends on both a coding rate and the channel coding scheme used. To satisfy the target bit-flip probability, ${\sf BER}(\{i\},n_{i})$ must not exceed $\mu_{i}$. Under these considerations, we formulate the bit-level UEP problem to minimize the total blocklength as follows:
${\textbf{(P1)}}\quad\min_{\{n_{i}\}_{i=1}^{K}}$~ $\sum_{i=1}^{K}n_{i}$ (9a)
s.t. ${\sf BER}(\{i\},n_{i})\leq\mu_{i},~\forall i,$ (9b)
$n_{i}\in\mathbb{Z}^{+},~\forall i.$ (9c)
III-B Channel Coding for Bit-Level UEP
As a simple yet practical approach to solving the bit-level UEP problem (P1), we apply repetition coding to each semantic bit using bounded distance decoding. The bit error probability under bounded distance decoding with $R$ repetitions for the $i$-th semantic bit is given by
${\sf BER}(\{i\},R)={\sf BER}_{\rm rep}(R)\triangleq\sum_{j=\lceil R/2\rceil}^{R}\binom{R}{j}\epsilon^{j}(1-\epsilon)^{R-j},$ (10)
for all $i$, where ${\sf BER}_{\rm rep}(R)$ is the error probability of $R$-repetition code, and $\epsilon$ is a bit-flip probability of each coded bit. For simplicity, we assume that each semantic bit is modulated using binary phase-shift keying (BPSK) with equal power $P_{\rm trans}$. Then, the bit-flip probability of each coded bit is given by $\epsilon=Q(\sqrt{2P_{\rm trans}/\sigma^{2}})$. It is worth noting that the above formulation is applicable to various modulation and power allocation schemes, provided that the value of $\epsilon$ is appropriately determined according to the specific modulation and power allocation strategy.
Since an $R$-repetition code can correct up to $\lfloor(R-1)/2\rfloor$ symbol errors, we restrict $R$ to be an odd number, which results in a perfect code. In this case, the problem (P1) can be decomposed into $K$ independent subproblems, each seeking the optimal number of repetitions required for a semantic bit to meet its target bit-flip probability. Accordingly, we focus on minimizing the repetition number $R$ given a target bit-flip probability $\mu_{i}$. This minimization problem can be expressed as
${(\textbf{P1}^{\prime})}\quad$ $\min_{r_{i}}\quad R_{i}$ (11a)
$~~\text{s.t.}\quad{\sf BER}_{\rm rep}(R_{i})\leq\mu_{i},$ (11b)
[…]
where the constraints (11c) and (11d) enforce the repetition number to be an odd number.
Since ${\sf BER}_{\rm rep}(2r+1)$ is strictly decreasing in $r$, the minimum value of $r\in\mathbb{Z}^{+}$ which satisfies ${\sf BER}_{\rm rep}(2r+1)\leq\mu$ can be determined using a bisection method. To establish the initial points for the bisection method, we set an upper bound of $r$ as
$r^{\rm(ub)}=\left\lfloor\frac{1}{2}\left(\frac{\ln\mu}{\ln(2\sqrt{(\epsilon(1-\epsilon)})}-1\right)\right\rfloor.$ (12)
The proof of the boundedness of $r^{\rm(ub)}$ is given in the following proposition:
Proposition 1:
Suppose that $r^{\rm(ub)}$ is defined as in (12), given a target bit-flip probability $\mu$ and a coded-bit-flip probability $\epsilon$. Then, the following inequality holds:
${\sf BER}_{\rm rep}(2r^{\rm(ub)}+1)\leq\mu.$ (13)
Proof:
See Appendix A. ∎
Additionally, to account for the fact that $P_{\rm rep}(2r+1)$ can only be evaluated at integer values, we modify the bisection method to track both the ceiling and floor values of the midpoint during the search. The detailed procedure is described in Algorithm 1.
Algorithm 1 Bisection method for determining the required repetition number
[…]
2:Initialize: $r^{\rm(lb)}\leftarrow 0,r^{\rm(ub)}\leftarrow\left\lfloor\frac{1}{2}\left(\frac{\ln\mu_{i}}{\ln(2\sqrt{(\epsilon(1-\epsilon)})}-1\right)\right\rfloor$
3:if $\epsilon<\mu_{i}$ then
4:   $r_{i}\leftarrow 0$
5:   Break
6:else
7:   repeat
8:   $c^{\rm(lb)}\leftarrow\left\lfloor\frac{r^{\rm(lb)}+r^{\rm(ub)}}{2}\right\rfloor,c^{\rm(ub)}\leftarrow\left\lceil\frac{r^{\rm(lb)}+r^{\rm(ub)}}{2}\right\rceil$
9:   if $P_{\rm rep}(2c^{\rm(ub)}+1)>\mu_{i}$ then
10:    $r^{\rm(lb)}\leftarrow c^{\rm(ub)}$
11:   else if $P_{\rm rep}(2c^{\rm(lb)}+1)<\mu_{i}$ then
12:    $r^{\rm(ub)}\leftarrow c^{\rm(lb)}$
13:   else
14:    $r_{i}\leftarrow c^{\rm(ub)}$
15:    Break
16:   end
17:   until
18:end
[…]
Despite the computational efficiency of the bisection method, applying it independently to each bit incurs excessive complexity proportional to $K$. To mitigate this complexity in determining the repetition numbers for all semantic bits, we develop a low-complexity approach for computing $\{R_{i}\}_{i=1}^{K}$. In this approach, we first sort [^2] In practice, this sorting necessitates corresponding reordering and inverse reordering at the transmitter and receiver, respectively, as illustrated in Fig. 1. This operation incurs no additional communication overhead since the learned bit-flip probabilities are shared between the transmitter and receiver during offline training, making this assumption practical. the semantic bits $\bm{b}$ in ascending order based on their target bit-flip probabilities $\mu_{i}$}. Then, we leverage the ordering of the target bit-flip probabilities $\{\mu_{i}\}_{i=1}^{K}$, which induces a monotonic constraint on $\{R_{i}\}_{i=1}^{K}$, i.e., $R_{i}\geq R_{i+1}$ for all $i$. Specifically, we determine the maximum repetition number, $R_{1}$, by applying the bisection method to the semantic bit with the most stringent requirement, $\mu_{1}$. Subsequently, we identify the transition points at which the repetition number changes, which occur when
$\mu_{i}<\mu_{i+1}\Rightarrow R_{i}>R_{i+1}.$ (14)
This approach requires only a single invocation of the bisection method, thereby significantly reducing the computational complexity compared to the naïve per-bit optimization strategy given in Algorithm 1. The detailed procedure is described in Algorithm 2.
Algorithm 2 Low complexity algorithm for determining the required repetition numbers
[…]
4:for $i=1$ to $K$ do
5:   if $\mu_{i}<\mu^{\rm(tr)}$ then
6:   $R_{i}\leftarrow R$
7:   else
8:   $\mu^{\rm(tr)}\leftarrow P_{\rm rep}(R-2)$
9:   $R\leftarrow R-2$
10:   $R_{i}\leftarrow R$
11:   end
12:   if $R=1$ then
13:   $R_{j}=1$$\forall j\geq i$
14:   Break
15:   end
16:end
[…]
IV Channel Coding for Block-Level UEP
in Semantic Communication
Although repetition coding offers analytical tractability for satisfying the bit-level UEP constraint, it provides no channel coding gain and exhibits poor transmission efficiency compared to modern channel codes. In semantic communication based on the method in Sec. II-B, certain semantic bits may share similar target bit-flip probabilities. This observation suggests that applying block coding to such groups of bits can yield improved coding gain without significantly violating the UEP constraint. Motivated by this insight, we propose a block-level UEP framework that enables grouping of bits with similar target bit-flip probabilities to exploit block coding gains. The overall structure of the proposed framework is illustrated in Fig. 4.
IV-A Problem Formulation for Block-Level UEP
In the proposed block-level UEP framework, we partition the $K$ semantic bits into $G$ groups based on their target bit-flip probabilities. Let $\mathcal{K}_{i}$ denote the set of indices corresponding to the $i$-th group, consisting of $k_{i}$ bits, given by
$\mathcal{K}_{i}=\{\bar{k}_{i-1}+1,\bar{k}_{i-1}+2,\ldots,\bar{k}_{i-1}+k_{i}\},$ (15)
where $\bar{k}_{i-1}=\sum_{j=1}^{i-1}k_{j}$. We then jointly encode the semantic bits within each group using a block channel encoder, without explicitly applying the UEP principle. Instead, the encoder is designed to satisfy the most stringent protection level within the group, i.e., to support the minimum target bit-flip probability among the bits in the group. Specifically, an $(n_{i},k_{i})$ block code is applied to the semantic bits in group $\mathcal{K}_{i}$ such that the resulting bit error probability, denoted by $P_{\rm ib}(\mathcal{K}_{i},n_{i})$, satisfies the following inequality:
$P_{\rm ib}(\mathcal{K}_{i},n_{i})\leq\min_{j\in\mathcal{K}_{i}}\;\mu_{j},\;\forall i\in[G].$ (16)
Note that the allowable values of $(n_{i},k_{i})$ are constrained by the structure of the specific coding scheme [^3] For example, when employing a binary primitive BCH code with error-correcting capability 2 and codeword length $n_{i}\leq 255$, the valid $(n_{i},k_{i})$ pairs are given by $\{(15,7),(31,21),(63,51),(127,113),(255,239)\}$ where $n_{i}$ is determined by $k_{i}$ [25]..
Figure 4: An illustration of the channel encoding process in the proposed block-wise UEP framework.
Considering the constraints on bit grouping and coding rate, we formulate the block-level UEP problem as follows:
${\textbf{(P2)}}~~\min_{\{\mathcal{K}_{i},n_{i}\}_{i=1}^{G},G}$~~ $\sum_{i=1}^{G}n_{i}$ (17a)
s.t. $P_{\rm ib}(\mathcal{K}_{i},n_{i})\leq\min_{j\in K_{i}}\mu_{j},\;\forall i\in[G],$ (17b)
$\mathcal{K}_{i}\cap\mathcal{K}_{i^{\prime}}=\emptyset,\;\forall i\neq i^{\prime},$ (17c)
$\cup_{i}\mathcal{K}_{i}=[K],$ (17d)
$|\mathcal{K}_{i}|\in\overline{\mathcal{K}},\;\forall i\in[G],$ (17e)
$n_{i}\in\mathcal{N}(|\mathcal{K}_{i}|),\;\forall i\in[G],$ (17f)
$G\in\mathbb{Z}^{+},$ (17g)
where $\overline{\mathcal{K}}$ is the set of possible group sizes before encoding, and $\mathcal{N}(|\mathcal{K}_{i}|)$ is the set of possible group sizes after encoding given $|\mathcal{K}_{i}|$. In the above problem, the constraint (17c) makes all groups disjoint from each other, and (17d) makes the groups cover all bits. Note that the constraints on possible group sizes, $\mathcal{K}$ and $\mathcal{N}(|\mathcal{K}_{i}|)$, depend on the channel coding scheme applied.
IV-B Bit Grouping Algorithm
As seen in the block-level UEP problem (P2), proper grouping of the semantic bits is critical in maximizing the performance of the block coding approach. For example, when bits have similar target bit-flip probabilities, forming larger groups can reduce the overall blocklength and improve coding efficiency. In contrast, when the target probabilities differ significantly, using smaller groups allows for more precise adherence to individual protection requirements, thereby enhancing overall UEP performance. We formalize this intuition by using the normal approximation of the finite blocklength capacity, which approximates the rate $k/n$ of an $(n,k)$ code as
$\frac{k}{n}\approx\log_{2}(1+\textsf{SNR})-\sqrt{\frac{V}{n}}\frac{Q^{-1}(\textsf{BLER})}{\ln 2},$ (18)
where BLER is the block error rate (BLER) of an $(n,k)$ code and $V=1-(1+\textsf{SNR})^{-2}$ is the channel dispersion [26]. Based on this, we approximate a BLER expression for an $(n,k)$ code as follows:
$\textsf{BLER}(n,k)=Q(f(n,k)),$ (19)
where $f(n,k)=\ln 2\sqrt{\frac{n}{V}}(\log_{2}(1+\textsf{SNR})-\frac{k}{n})$ [27].
On the basis of the BLER expression given in (19), we prove that a blocklength can be reduced by grouping semantic bits with the same protection level. This result is formally stated in the following theorem:
Proposition 2:
Consider two block codes: an $(n_{1},k_{1})$ code and an $(n_{2},k_{2})$ code, whose BLERs are given as in (19). If $\textsf{BLER}(n_{1},k_{1})=\textsf{BLER}(n_{2},k_{2})$, there exists a positive integer $n_{3}<n_{1}+n_{2}$ such that
$\textsf{BLER}(n_{3},k_{1}+k_{2})=\textsf{BLER}(n_{1},k_{1}).$ (20)
Proof:
See Appendix B. ∎
Additionally, we formally characterize the condition under which grouping semantic bits results in a longer blocklength. This finding is stated in the following theorem:
Theorem 1:
Consider three block codes: an $(n_{1},k_{1})$ code, an $(n_{2},k_{2})$ code, and an $(n_{3},k_{1}+k_{2})$ code, whose BLERs are given as in (19). If $\textsf{BLER}(n_{3},k_{1}+k_{2})=\textsf{BLER}(n_{1},k_{1})$, then a necessary and sufficient condition for $n_{3}>n_{1}+n_{2}$ is
$\frac{Q^{-1}(\textsf{BLER}(n_{1},k_{1}))}{Q^{-1}(\textsf{BLER}(n_{2},k_{2}))}\geq\gamma_{\rm th}(n_{1},n_{2},k_{1},k_{2}),$ (21)
provided that $\gamma_{\rm th}(n_{1},n_{2},k_{1},k_{2})$ is given by
$\gamma_{\rm th}(n_{1},n_{2},k_{1},k_{2})$
$=\frac{\sqrt{2Y(\sqrt{Y^{2}+4(k_{1}+k_{2})C}-\sqrt{Y^{2}+4k_{1}C})+4k_{2}C}}{\sqrt{Y^{2}+4(k_{1}+k_{2})C}-\sqrt{Y^{2}+4k_{1}C}},$ (22)
where $C=\log_{2}(1+\textsf{SNR})$, and $Y=\sqrt{n_{1}}C-k_{1}/\sqrt{n_{1}}$.
Proof:
See Appendix C. ∎
Proposition 2 establishes the advantage of grouping semantic bits whose target protection levels are similar. Conversely, Theorem 1 emphasizes the necessity of splitting groups when semantic bits have significantly different protection levels. The left-hand side of (21) quantifies the BLER difference between two groups. Specifically, as $\textsf{BLER}(n_{2},k_{2})$ increases, the ratio $Q^{-1}(\textsf{BLER}(n_{1},k_{1}))/Q^{-1}(\textsf{BLER}(n_{2},k_{2}))$ also increases, making separation of the groups preferable. This intuitive result provides a quantitative criterion for deciding when to merge or separate groups. Motivated by this observation, we propose a grouping algorithm that minimizes the overall blocklength while ensuring that the target bit-flip probabilities are satisfied.
Proposition 2 indicates that grouping bits with comparable error protection requirements reduces blocklength, but it does not explicitly define what constitutes comparable. To address this, we leverage a repetition-based UEP criterion: each semantic bit $i$ is assigned a repetition number $R_{i}$ such that its bit-flip probability remains below the target $\mu_{i}$. Bits sharing the same repetition number inherently possess equivalent protection levels, as they receive identical error protection. We thus group bits according to identical $R_{i}$ values, exploiting coding gain without violating repetition-based UEP constraints. Specifically, the first group includes bits whose repetition numbers are $R_{\max}$, the second group includes bits with repetition number $R_{\max}-2$, and so forth, until the final group contains bits with the minimum repetition number (i.e., $R_{i}=3$). It is worth noting that block channel coding is not applied to semantic bits with a repetition number of one (i.e., $R_{i}=1$), as these bits already meet their target error protection requirements without the need for additional coding.
After this initial grouping, the total number of the non-singleton bit groups is $(\max_{i}R_{i}-1)/2$. To further reduce the overall blocklength, we utilize the criterion derived from Theorem 1. Although Theorem 1 originally provides an equivalent condition favoring separation, it implicitly produces the following condition for merging groups:
$\frac{Q^{-1}(\textsf{BLER}(n_{1},k_{1}))}{Q^{-1}(\textsf{BLER}(n_{2},k_{2}))}<\gamma_{\rm th}(n_{1},n_{2},k_{1},k_{2}).$ (23)
However, since the exact blocklengths $n_{1}$ and $n_{2}$ required to achieve the target bit-flip probabilities for groups of sizes $k_{1}$ and $k_{2}$ respectively are unknown, we approximate the BLER of any group $\mathcal{A}$ as follows:
$\textsf{BLER}(\mathcal{A})\approx 1-\Big{(}1-\min_{i\in\mathcal{A}}\mu_{i}\Big{)}^{|\mathcal{A}|},$ (24)
which estimates the probability that at least one bit within a group of size $|\mathcal{A}|$ is erroneous [28]. Using this approximation, we further obtain
$Y\approx\frac{\sqrt{V}}{\ln 2}Q^{-1}(\textsf{BLER}(\mathcal{A})),$ (25)
allowing us to apply the merging condition effectively. The proposed grouping algorithm is detailed in Algorithm 3.
Algorithm 3 Proposed bit grouping algorithm
[…]
3:for $p=1$ to $(R_{\max}-1)/2$ do
4:   $R\leftarrow R_{\max}-2(p-1)$
5:   $\mathcal{J}_{p}\leftarrow\{j:R_{j}=R\}$
6:end
[…]
10:for $p=1$ to $(R_{\max}-1)/2-1$ do
11:   $\mathcal{B}\leftarrow\mathcal{J}_{p+1}$
12:   $\mu^{(1)}\leftarrow\min_{i\in\mathcal{A}}\mu_{i}$
13:   $\mu^{(2)}\leftarrow\min_{i\in\mathcal{B}}\mu_{i}$
14:   $\textsf{BLER}^{(1)}\leftarrow 1-(1-\mu^{(1)})^{|\mathcal{A}|}$
15:   $\textsf{BLER}^{(2)}\leftarrow 1-(1-\mu^{(2)})^{|\mathcal{B}|}$
16:   $Y\leftarrow\frac{\sqrt{V}}{\ln 2}Q^{-1}(\textsf{BLER}^{(1)})$
[…]
18:   if $Q^{-1}(\textsf{BLER}^{(1)})/Q^{-1}(\textsf{BLER}^{(2)})<\gamma_{\rm th}$ then
19:   $\mathcal{A}\leftarrow\mathcal{A}\cup\mathcal{B}$
20:   $\mathcal{K}_{g}\leftarrow\mathcal{A}$
21:   else
22:   $\mathcal{K}_{g}\leftarrow\mathcal{A}$, $\mathcal{K}_{g+1}\leftarrow\mathcal{B}$
23:   $g\leftarrow g+1$
24:   $\mathcal{A}\leftarrow\mathcal{J}_{i+1}$
25:   end
26:end
[…]
Algorithm 4 Algorithm for fitting group sizes
[…]
4:while $k_{\rm acc}<K_{\rm group}$ do
5:   $k^{\star}\leftarrow\arg\min_{k\in\overline{\mathcal{K}}}||\mathcal{K}_{g}|-k|$
6:   if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}\leq K_{\rm group}$ then
7:   while $|\mathcal{K}_{g}|<k^{\star}$ do
8:    $\mathcal{K}_{g+1}\leftarrow\mathcal{K}_{g}\cup\mathcal{K}_{g+1}$
9:    $g\leftarrow g+1$
10:   end
11:   end
12:   if $|\mathcal{K}_{g}|\geq k^{\star}$ then
13:   $\Delta k\leftarrow|\mathcal{K}_{g}|-k^{\star}$
14:   $\Delta\mathcal{K}\leftarrow\mathcal{K}_{g}[-\Delta k:\text{end}]$
15:   $\mathcal{K}_{i}^{\prime}\leftarrow\mathcal{K}_{g}\setminus\Delta\mathcal{K}$
16:   $\mathcal{K}_{g+1}\leftarrow\mathcal{K}_{g+1}\cup\Delta\mathcal{K}$
17:   $k_{\rm acc}\leftarrow k_{\rm acc}+k^{\star}$
18:   else if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}>K_{\rm group}$ then
19:   if $k_{\rm acc}+k^{\star}>K$ then
20:    $\mathcal{K}_{i}^{\prime}\leftarrow\{\mathcal{K}_{g}[1],\mathcal{K}_{g}[2],\ldots,K,0,\ldots,0\}$ so that $|K_{i}^{\prime}|=k^{\star}$
21:    $\mathcal{J}_{\rm single}^{\prime}\leftarrow\emptyset$
22:   else
23:    $\mathcal{K}_{i}^{\prime}\leftarrow\{\mathcal{K}_{g}[1],\mathcal{K}_{g}[2],\ldots,k_{\rm acc}+k^{\star}\}$ so that $|K_{i}^{\prime}|=k^{\star}$
24:    $\mathcal{J}_{\rm single}^{\prime}\leftarrow\{k_{\rm acc}+k^{\star}+1,k_{\rm acc}+k^{\star}+2,\ldots,K\}$
25:   end
26:   $k_{\rm acc}\leftarrow k_{\rm acc}+k^{\star}$
27:   end
28:   $i\leftarrow i+1$, $g\leftarrow g+1$
29:end
[…]
IV-C Channel Coding for Block-Level UEP
Algorithm 5 Algorithm for block-level UEP
[…]
4:for $g=1$ to $G^{\prime}$ do
5:   $r_{g}\leftarrow\max\;r\in\mathcal{R}$ s.t. $T(|\mathcal{K}_{g}^{\prime}|,r)<\min_{i\in\mathcal{K}_{g}^{\prime}}\mu_{i}$
6:end
[…]
After determining the $G$ groups $\{\mathcal{K}_{g}\}_{g=1}^{G}$, the semantic bits within each group are jointly encoded by a channel encoder designed to support the minimum target bit-flip probability among the bits in that group. To utilize the capacity-based analysis in Proposition 2 and Theorem 1, which serve as theoretical guidelines under the assumption of capacity-achieving codes, we employ modern channel coding schemes to minimize the gap between theory and practice. Specifically, we employ polar codes and LDPC codes. Polar codes are known to achieve capacity at infinite blocklength and outperforming LDPC in shorter lengths [29], whereas LDPC codes typically exhibit better performance at moderate to high code rates and longer blocklength [30]. Therefore, we apply polar codes and LDPC codes for short and large blocklengths, respectively. Further details on adopted channel coding schemes are provided in Sec. V.
To ensure compatibility with practical polar/LDPC codes and reduce implementation complexity, we constrain the possible group sizes to a predefined set $\overline{\mathcal{K}}$. Note that larger group sizes tend to incur higher decoding complexity due to increased dimensions of the parity-check matrix and the iteration overhead in belief propagation decoding. Furthermore, to maintain tractable memory complexity, we restrict the possible code rates to the set $\mathcal{R}$, e.g. $\mathcal{R}=\{1/2,1/3,1/4\}$, motivated by modulation and coding scheme (MCS) tables commonly found in modern wireless standards [31]. Consequently, we modify the group sizes produced by Algorithm 3 to align with the set $\overline{\mathcal{K}}$. This adjustment process, detailed in Algorithm 4, merges or splits adjacent groups as necessary, ensuring each resulting group $\mathcal{K}_{i}^{\prime}$ closely matches an element of $\overline{\mathcal{K}}$. Specifically, if a group size is smaller than the required code length, subsequent groups are merged to reach the necessary size. Conversely, if a group is excessively large, surplus bits are moved to the following group. This ensures compatibility with practical polar and LDPC codebooks.
To summarize the complete block-level UEP procedure, we present a unified algorithm incorporating the previously discussed grouping and rate-selection steps. As detailed in Algorithm 5, the procedure begins by executing the grouping algorithm (Algorithm 3), followed by adjustments to the group sizes to conform to the allowed group sizes (Algorithm 4). Subsequently, for each group, the most suitable code rate is selected from the predefined set $\mathcal{R}$ by referencing a lookup table $T(k,r)$ that maps group sizes and code rates to achievable bit-flip probabilities. The selected code rate ensures that the bit-flip probability of all bits within the group does not exceed the smallest target bit-flip probability among its members. This integrated approach enables the block-level UEP framework to achieve varying protection levels systematically, while remaining fully compatible with practical codebooks.
V Simulation Results
In this section, we evaluate the superiority of the proposed UEP frameworks for an image transmission task with the MNIST, CIFAR-10, and CIFAR-100 datasets [32, 33]. The MNIST dataset contains 70,000 grayscale images of handwritten digits of size 28x28 pixels, divided into 60,000 training and 10,000 test samples. The CIFAR-10 and CIFAR-100 datasets consist of 60,000 color images of size 32×32×3 pixels. CIFAR-10 is split into 50,000 training and 10,000 test samples across 10 classes, while CIFAR-100 includes 100 classes with the same train/test split.
TABLE I: Semantic encoder and decoder architectures for the MNIST dataset.
Index Layer Name NN Units Activation
Encoder
1 CONV1 32, 3 $\times$ 3 PReLU
2 CONV2 64, 3 $\times$ 3 PReLU
3 CONV3 64, 5 $\times$ 5 PReLU
4 CONV4 8, 5 $\times$ 5 None
5 Flatten - None
Decoder
6 Unflatten - None
7 DECONV1 64, 5 $\times$ 5 PReLU
8 DECONV2 64, 5 $\times$ 5 PReLU
9 DECONV3 32, 3 $\times$ 3 PReLU
10 DECONV4 1, 4 $\times$ 4 None
TABLE II: Semantic encoder and decoder architectures for the CIFAR-10 and CIFAR-100 datasets.
Index Layer Name NN Units Activation
Encoder
1 CONV1 64, 5 $\times$ 5 PReLU
2 CONV2 128, 5 $\times$ 5 PReLU
3 CONV3 128, 5 $\times$ 5 PReLU
4 CONV4 128, 5 $\times$ 5 PReLU
5 CONV5 24, 5 $\times$ 5 None
6 Flatten - None
Decoder
7 Unflatten - None
8 DECONV1 128, 5 $\times$ 5 PReLU
9 DECONV2 128, 5 $\times$ 5 PReLU
10 DECONV3 128, 5 $\times$ 5 PReLU
11 DECONV4 64, 5 $\times$ 5 PReLU
12 DECONV5 3, 5 $\times$ 5 None
The encoder–decoder pair and the bit-flip probabilities $\{\mu_{i}\}$ are jointly trained as defined in (8). We employ convolutional neural network (CNN)-based autoencoder architectures, following the design in [6], which are detailed in Table I and Table II. The feature vector is quantized using 8 bits. The symbol transmission power is set to $P_{\rm trans}=0$ dBW, and the SNR is fixed at 0 dB under an AWGN channel.
In the simulations, we compare the following channel coding frameworks:
  • Bit-UEP: This is the proposed bit-level UEP framework. Based on the learned target bit-flip probabilities $\{\mu_{i}\}$, we apply Algorithm 2 to assign an individual repetition number to each bit.
  • Block-UEP: This is the proposed block-level UEP framework. Using the learned $\{\mu_{i}\}$, we apply Algorithm 5 to group bits with similar protection requirements. The set of allowable group sizes is $\mathcal{K}=\{128,\,256,\,512,\,1024\}$. Each group is encoded using either a polar code (for group sizes of 128 and 256) or an LDPC code (for group sizes of 512 and 1024), where the code rate is selected from a lookup table with available code rates $\mathcal{R}=\{3/4,\,2/3,\,15/24,\,14/24,\,13/24,\,1/2,\,1/3\}$. Both polar and LDPC codes are implemented according to the 5G standard: polar codes utilize CRC-aided list decoding with a list size of 8, while LDPC codes are constructed using the standard’s two base graphs and a lifting procedure [31].
  • Repetition ($R_{\rm fix}$): All bits are encoded using the repetition code with a fixed repetition number $R_{\rm fix}$.
  • LDPC: This is an equal error protection scheme using an LDPC code. Bit grouping is performed identically to the Block-UEP scheme. However, all groups are encoded using LDPC codes with a fixed rate, which is selected to ensure that the total blocklength matches that of the Block-UEP scheme.
  • Polar: This is an equal error protection scheme using a polar code. Since a large group size is prohibitive due to polar code’s implementation complexity, all semantic bit groups are uniformly set to size 256. All groups use the same code rate selected to ensure that the total blocklength matches that of the Block-UEP scheme.
  • Genie: This ideal baseline assumes perfect transmission with no bit-flips, i.e., all bits are received without error.
Figure 5: Comparison of PSNR versus total blocklength for various channel coding frameworks using the CIFAR-10 dataset with $\lambda=10^{-4}$ ($K=12288$).
Figure 6: Comparison of PSNR versus total blocklength for various channel coding frameworks using the CIFAR-10 dataset with $\lambda=10^{-6}$ ($K=12288$).
In Figs. 5 and 6, we compare the PSNR and corresponding total blocklength for various channel coding frameworks, using the CIFAR-10 dataset with $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results demonstrate that the proposed bit-level UEP framework (Bit-UEP) consistently achieves higher PSNR compared to fixed-rate repetition coding (Repetition) under both training settings. Specifically, the fixed-rate scheme with $R_{\rm fix}=8$ consumes slightly more blocklength than Bit-UEP, yet suffers from a significantly lower PSNR. On the other hand, the fixed-rate scheme with $R_{\rm fix}=12$ achieves comparable PSNR to Bit-UEP but at the cost of significantly larger blocklength usage. Moreover, coding gain is evident when comparing Bit-UEP and Block-UEP. Block-UEP achieves higher PSNR with even less blocklength usage compared to Bit-UEP. Additionally, Block-UEP outperforms traditional equal rate schemes (LDPC and Polar) in terms of PSNR while maintaining the same blocklength across both datasets. This result demonstrates the importance and effectiveness of dynamic grouping of the semantic bits according to their protection requirements.
Similarly, in Figs. 7 and 8, we compare the PSNR and total blocklegnth across various channel coding frameworks using the CIFAR-100 dataset, with neural networks trained under $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results again confirm the superiority of unequal error protection over traditional equal protection schemes. From the consistent observations on both CIFAR-10 and CIFAR-100, we conclude that the proposed UEP frameworks provide clear advantages for semantic communication in terms of both reconstruction quality and transmission efficiency. Notably, the performance of Block-UEP approaches the genie-aided upper bound in both datasets, indicating its effectiveness in achieving near-optimal reconstruction quality with an improved transmission efficiency. In Fig. 9, we visualize this effect with reconstructed CIFAR-100 images. The figure demonstrates that Block-UEP offers the highest image quality, while the equal error protection schemes (LDPC and Polar) suffer from image quality degradation.
Figure 7: Comparison of PSNR versus total blocklength for various channel coding frameworks using the CIFAR-100 dataset with $\lambda=10^{-4}$ ($K=12288$).
Figure 8: Comparison of PSNR versus total blocklength for various channel coding frameworks using the CIFAR-100 dataset with $\lambda=10^{-6}$ ($K=12288$).
Figure 9: Visualization of reconstructed images and corresponding PSNR and SSIM values for various channel coding frameworks using CIFAR-100 dataset with $\lambda=10^{-4}$.
TABLE III: PSNR of the proposed block-level UEP framework with different orderings of semantic bits.
Dataset($\lambda$) Proposed Random Reverse
CIFAR-10($10^{-6}$) 42.24 29.33 20.00
CIFAR-10($10^{-4}$) 42.19 20.69 11.44
MNIST($10^{-6}$) 43.80 26.06 13.81
MNIST($10^{-4}$) 43.45 18.02 9.40
In Table III, we compare the PSNR of the proposed block-level UEP framework with different orderings of semantic bits. In the proposed ordering, the semantic bits are sorted in ascending order of their target bit-flip probabilities, as described in Sec. III and Sec. IV, and Algorithm 5 is applied to this sorted sequence. For comparison, we consider two alternative orderings: random and reverse. As shown in Table III, the proposed ordering achieves the highest PSNR, while the reverse ordering yields the worst performance. This result confirms that bits with lower target bit-flip probabilities are more critical for image reconstruction. In Fig. 10, we also visualize this effect with reconstructed CIFAR-10 images. The figure demonstrates that reversing the bit order leads to severe degradation in image quality.
In Figs. 11 and 12, we compare the PSNR and total blocklength of the proposed block-level UEP framework with different values of $\lambda$ using the MNIST and CIFAR-10 datasets, respectively. The regularization parameter $\lambda$ penalizes the assignment of excessively low bit-flip probabilities to task-irrelevant bits in the loss function (8). As shown in the figures, a larger $\lambda$ tends to reduce PSNR, as it limits the model’s capacity to minimize reconstruction error. However, it also significantly decreases total blocklength in both datasets. These results suggest that selecting an appropriate value of $\lambda$ allows for a trade-off between reconstruction quality and transmission efficiency, depending on the desired task performance level.
Figure 10: Visualization of reconstructed images and corresponding PSNR and SSIM values of the proposed block-level UEP framework with different orderings of semantic bits for $\lambda=10^{-4}$.
Figure 11: PSNR and total blocklength the proposed block-level UEP framework with various $\lambda$ values using the MNIST dataset ($K=3136$).
Figure 12: PSNR and total blocklength the proposed block-level UEP framework with various $\lambda$ values using the CIFAR-10 dataset ($K=12288$).
VI Conclusion
In this paper, we explored the UEP problem in digital semantic communication by leveraging learned bit-flip probabilities as target protection levels. We proposed two coding frameworks: a bit-level UEP strategy based on repetition coding and a block-level UEP approach using polar codes and LDPC codes. The bit-level framework minimizes total blocklength while ensuring per-bit reliability, whereas the block-level framework enhances coding efficiency by grouping bits with similar protection requirements. For the block-level framework, we introduced a bit-grouping algorithm guided by finite blocklength capacity analysis. Simulation results demonstrated that both frameworks significantly improve task performance and transmission efficiency compared to conventional methods. These results highlight the importance of bit-level error protection in semantic communication and lay the groundwork for future research on resource-efficient and reliability-aware semantic transmission strategies.
A promising avenue for future research is to extend the proposed UEP framework to multi-user semantic communication scenarios, where users may experience heterogeneous channel conditions and demand different levels of quality-of-service (QoS). In such settings, efficient multiple access, user grouping and adaptive protection strategies become critical. Another important direction is to incorporate multi-antenna techniques to further improve transmission efficiency.
Appendix A Proof of Proposition 1
To show the boundedness by $r^{\rm(ub)}$, define a random variable $X$ which follows the binomial distribution ${\rm Bino}(R,\,\epsilon)$ so that $P_{\rm rep}(R)={\rm Pr}(X\geq R/2)$. Now, suppose that $\epsilon<1/2$ without loss of generality, and define $\delta=1/(2\epsilon)-1$. Then, $P_{\rm rep}(R)$ is upper bounded as follows:
$P_{\rm rep}(R)$ $={\rm Pr}\left(X\geq\frac{R}{2}\right)$ (26)
$={\rm Pr}(X\geq(1+\delta)\epsilon R)$ (27)
$\overset{(a)}{\leq}\exp(-R\cdot{\rm KL}((1+\delta)\epsilon||\epsilon))$ (28)
$=\exp\left(-R\cdot{\rm KL}\left(\frac{1}{2}||\epsilon\right)\right)$ (29)
$=\left(2\sqrt{\epsilon(1-\epsilon)}\right)^{R},$ (30)
where the inequality (a) is followed by Chernoff bound with Kullback–Leibler divergence [34].
Therefore, the definition of $r^{\rm(ub)}$ gives
 $\;2r^{\rm(ub)}+1\leq\frac{\ln\mu}{\ln(2\sqrt{\epsilon(1-\epsilon)})}$ (31)
$\Rightarrow\ln P_{\rm rep}(2r^{\rm(ub)}+1)\leq(2r^{\rm(ub)}+1)\ln(2\sqrt{\epsilon(1-\epsilon)})$ (32)
$\Rightarrow P_{\rm rep}(2r^{\rm(ub)}+1)\leq\mu.$ (33)
Appendix B Proof of Proposition 2
From $\textsf{BLER}(n_{1},k_{1})=\textsf{BLER}(n_{2},k_{2})$, we have
$Y\triangleq\sqrt{n_{1}}C-\frac{k_{1}}{\sqrt{n_{1}}}=\sqrt{n_{1}}C-\frac{k_{1}}{\sqrt{n_{1}}},$ (34)
which gives $k_{1}=n_{1}C-Y\sqrt{n_{1}}$ and $k_{2}=n_{2}C-Y\sqrt{n_{2}}$, where $C=\log_{2}(1+\textsf{SNR})$. By plugging this into $k_{1}+k_{2}$,
$\quad\sqrt{n_{1}+n_{2}}C-\frac{k_{1}+k_{2}}{\sqrt{n_{1}+n_{2}}}$ (35)
$=\sqrt{n_{1}+n_{2}}C-\frac{(n_{1}+n_{2})C-Y(\sqrt{n_{1}}+\sqrt{n_{2}})}{\sqrt{n_{1}+n_{2}}}$ (36)
$=Y\frac{\sqrt{n_{1}}+\sqrt{n_{2}}}{\sqrt{n_{1}+n_{2}}}>Y.$ (37)
Therefore,
$\quad\textsf{BLER}(n_{1}+n_{2},k_{1}+k_{2})$ (39)
$=Q\left(\frac{\ln 2}{\sqrt{V}}\left(\sqrt{n_{1}+n_{2}}R-\frac{k_{1}+k_{2}}{\sqrt{n_{1}+n_{2}}}\right)\right)$ (40)
$<Q\left(\frac{\ln 2}{\sqrt{V}}Y\right)$ (41)
$=\textsf{BLER}(n_{1},k_{1}).$ (42)
By the monotonicity of $\textsf{BLER}(n,k)$ in $n$, there exists $n_{3}$ such that $\textsf{BLER}(n_{3},k_{1}+k_{2})=\textsf{BLER}(n_{1},k_{1})$ and $n_{3}<n_{1}+n_{2}$.
Appendix C Proof of Theorem 1
From $\textsf{BLER}(n_{1},k_{1})=\textsf{BLER}(n_{3},k_{1}+k_{2})$, we have
$Y$ $=\sqrt{n_{1}}C-\frac{k_{1}}{\sqrt{n_{1}}}$ (43a)
$=\sqrt{n_{3}}C-\frac{k_{1}+k_{2}}{\sqrt{n_{3}}}.$ (43b)
Also, define $\gamma\triangleq Q^{-1}(\textsf{BLER}(n_{1},k_{1}))/Q^{-1}(\textsf{BLER}(n_{2},k_{2}))$ so that measure the difference of BLER between $(n_{1},k_{1})$ code and $(n_{2},k_{2})$ code, which gives
$Y=\gamma\left(\sqrt{n_{2}}C-\frac{k_{2}}{\sqrt{n_{2}}}\right).$ (44)
Since (43a), (44) and (43b) are all quadratic equation of $\sqrt{n_{1}}$, $\sqrt{n_{2}}$ and $\sqrt{n_{3}}$, respectively, quadratic formula gives
$\sqrt{n_{1}}$ $=\frac{1}{2C}\left(Y+\sqrt{Y^{2}+4k_{1}C}\right),$ (45)
$\sqrt{n_{2}}$ $=\frac{1}{2C}\left(\frac{Y}{\gamma}+\sqrt{\frac{Y^{2}}{\gamma^{2}}+4k_{2}C}\right),$ (46)
$\sqrt{n_{3}}$ $=\frac{1}{2C}\left(Y+\sqrt{Y^{2}+4(k_{1}+k_{2})C}\right).$ (47)
Therefore,
$n_{3}-(n_{1}+n_{2})=\frac{Y}{2C^{2}}\bigg{(}\sqrt{Y^{2}+4(k_{1}+k_{2})C}$
$\qquad\quad-\sqrt{Y^{2}+4k_{1}C}-\frac{1}{\gamma}\sqrt{\frac{Y^{2}}{\gamma^{2}}+4k_{2}C}-\frac{Y}{\gamma^{2}}\bigg{)}.$ (48)
From this relation, $n_{3}-(n_{1}+n_{2})>0$ is equivalent to
$A-\frac{Y}{\gamma^{2}}>\frac{1}{\gamma}\sqrt{\frac{Y^{2}}{\gamma^{2}}+4k_{2}C},$ (49)
where $A\triangleq\sqrt{Y^{2}+4(k_{1}+k_{2})C}-\sqrt{Y^{2}+4k_{1}C}$. This inequality holds if and only if
$A^{2}-\frac{2AY}{\gamma^{2}}-\frac{4k_{2}C}{\gamma^{2}}>0,$ (50)
which is equivalent to
$\gamma>\gamma_{\rm th}(n_{1},n_{2},k_{1},k_{2}).$ (51)
This condition implies that merging two bit groups whose BLERs sufficiently differ uses more blocklength.